OpenStudy (anonymous):

Find the domain f(x)= 2x/x^2-4x-5

OpenStudy (ranga):

Is the entire (x^2 - 4x - 5) in the denominator?

OpenStudy (anonymous):

yes

OpenStudy (ranga):

factor the denominator: x^2 - 4x - 5 = x^2 - 5x + x - 5 x(x-5) + 1(x-5) (x+1)(x-5) f(x) = 2x / { (x+1)(x-5) } The denominator cannot be zero. Therefore the domain should exclude what 2 values of x?

OpenStudy (anonymous):

-1 and 5